find an equetion of parabola with the axis is parallel to y-axis & the parabola passes through (-4,21),(-2,11),(4,5)
Answer by lwsshak3(11628)

find an equation of parabola with the axis is parallel to y-axis & the parabola passes through (-4,21),(-2,11),(4,5)
***
General form of equation for a parabola:
y=ax^2+bx+c
..
(-4,21), 21=16a-4b+c
(-2,11), 11=4a-2b+c
(4,5),.... 5=16a+4b+c
..
21=16a-4b+c
11=4a-2b+c
subtract to eliminate c
10=12a-2b
..
11=4a-2b+c
5=16a+4b+c
subtract to eliminate c
6=-12a-6b
..
10=12a-2b
6=-12a-6b
add to eliminate a
16=-8b
b=-2
12a=10+2b=10-4=6
a=6/12=1/2
c=11-4a+2b=11-2-4=5
..
a=1/2, b=-2, c=5
Equation: y=(1/2)x^2-2x+5